7. (10 points) Find the general solution to the homogeneous system of DE: -1 x = Ax where A -2 = [ 21
Answer #1

Ax = ا A : - | 2 - ) + 0 = کد 2 دا د ن : ) کے۔ ما 2 + ا- - 2 2- ا- دهه For d=-1+2i CA -d2 Il v2=0 LA-d, Ilv, = 0 2 V \ نات -ا